### v1

# Block structures of graphs and quantum isomorphism

Abstract.
We prove that for every pair of quantum isomorphic graphs, their block trees and their block graphs are isomorphic, and that such an isomorphism can be chosen so that the corresponding blocks are quantum isomorphic. As a corollary of this result, we obtain that a minimal pair of quantum isomorphic graphs which are not isomorphic consists of 2-connected graphs.
1Université Paris-Saclay, CNRS, Laboratoire de Mathématiques d’Orsay, 91405 Orsay, France
2KU Leuven, Department of Mathematics, Celestijnenlaan 200B box 2400, BE-3001 Leuven, research supported by the grant 11PAL24N funded by the Research Foundation Flanders (FWO)
3Chaire Combinatoire, Collège de France, Université PSL, 75005, Paris, France
1. Introduction
In recent years, the notion of quantum isomorphism of graphs gained a major place in the study of noncommutative properties of finite graphs. Indeed, this seemingly candid definition turned out to be related to deep mathematical questions, such as the study of $\mathcal{F}$-isomorphism [MR19], the representation of finitely presented groups [AMR+19, Slo20], or the study of quantum automorphism groups of graphs. While striking results allow for a global description of quantum isomorphism (see [MR19] for characterisations related to homomorphism count or representations of compact quantum groups), very little is explicitly known about it. For instance, we still do not know a smallest pair of quantum isomorphic graphs which are not isomorphic: the smallest examples known have 24 vertices [AMR+19], and in [Meu23] it is shown that such graphs need at least 5 vertices (see Section 6 for more open problems).
To attack such problems, as well as for explicit computations of quantum automorphism groups of graphs, it is of great importance to understand how quantum isomorphism behaves with respect to the structure of graphs. This understanding for the decomposition of a graph into its connected components, which was partially used in several works [LMR20, DBKR+23], was only recently established in full generality [Meu23], where it is shown that if two graphs are quantum isomorphic, then their connected components are in bijection and two-by-two quantum isomorphic themselves.
Our main result is to go one step further and understand how the decomposition of a graph into its 2-connected components behaves under quantum isomorphism for two important block structures. The answer is the following (see Theorem 5.3).
Theorem A.
If $G$ and $H$ are two quantum isomorphic graphs, then
  1. (1)
    there exists $\alpha:\mathcal{T}(G)\rightarrow\mathcal{T}(H)$ such that
    • $\alpha$ is an isomorphism between the block trees of $G$ and $H$,
    • for every block $b$ of $G$, we have $b\simeq_{q}\alpha(b)$.
  2. (2)
    there exists $\beta:\mathcal{B}(G)\rightarrow\mathcal{B}(H)$ such that
    • $\beta$ is an isomorphism between the block graphs of $G$ and $H$,
    • for every block $b$ of $G$, we have $b\simeq_{q}\beta(b)$.
As an immediate corollary, we obtain that a minimal pair of isomorphic graphs which are not quantum isomorphic is 2-connected (see Corollary 5.4). This theorem will also be the key to a systematic study of quantum properties of block graphs as well as the computation of their quantum automorphism groups in a forthcoming paper.
Our strategy to prove this result is inductive. More precisely, we will define an operation $\Gamma$ which splits a given graph into a disjoint union of new graphs with the property that each connected component has less blocks that the original graph. Even though the construction is elementary, it is technical because we have to keep track of blocks and vertices along which the decomposition is made. To do this, we will introduce a notion of anchored graph and define the $\Gamma$ operation at this level. We will then in the end check that the results obtained for anchored graphs can be lifted back to the original graphs.
In Section 2, we recall the necessary definitions and introduce our notations. In Section 3, we state and prove the fundamental theorem for quantum isomorphism of partitioned graphs (Theorem 3.3), and we obtain that blocks are preserved by magic unitaries (Lemma 3.5). In Section 4, we introduce the operation $\Gamma$ to obtain our main technical result: if $G$ and $H$ are quantum isomorphic, then so are $\Gamma(G)$ and $\Gamma(H)$ (this is actually defined for anchored graphs, see Lemma 4.8 for the precise statement). Section 5 is devoted to the proof of Theorem A and other related results. Finally, Section 6 contains some open problems that we briefly mention here, and leave the proper definitions and details to the related section.
  1. (1)
    Does there exist a pair of quantum isomorphic graphs $G$ and $H$ whose connectivities are not the same?
  2. (2)
    If yes, what is the maximum $k$ such that $k$-connectivity is preserved under quantum isomorphism? That is, what is the maximum $k$ such that for every $k$-connected graph $G$ we have that every graph quantum isomorphic to $G$ is $k$-connected?
  3. (3)
    Does there exist an irreducible pair $(G,H)$ of quantum isomorphic graphs that are not quantum isomorphic and such that $G$ is not 3-connected?
  4. (4)
    Are there infinitely many irreducible pairs of quantum isomorphic graphs that are not isomorphic?
2. Preliminaries
Let us start by introducing some notation and providing the preliminary results.
Graph theoretical notions
For any graph theoretical notion not defined here, we refer to [BM08].
All graphs in this paper are finite and simple – that is, loopless and without multiple edges. We denote the vertex set and the edge set of a graph $G$ respectively by $V(G)$ and $E(G)$, and its adjacency matrix by $\operatorname{Adj}(G)$. We may write $xy$ for an element $\{x,y\}\in E(G)$. For $x\in V(G)$ the neighbourhood of $x$ is defined as $N_{G}(x)=\{y\in V(G)\mid xy\in E(G)\}$ and the closed neighbourhood of $x$ is defined as $N_{G}[x]=N_{G}(x)\cup\{x\}$.
A graph morphism (or simply a morphism) from a graph $G$ to a graph $H$ is a function $\phi:V(G)\to V(H)$ such that for all $\{x,y\}\in E(G)$ we have $\phi(x)\phi(y)\in E(H)$. The morphism $\phi$ is an isomorphism if and only if it is a bijection and for every $x,y\in V(G)$ such that $xy\notin E(G)$ we have $\phi(x)\phi(y)\notin E(H)$.
We use the notation $G[S]$ for the subgraph $H$ of $G$ induced by a set $S\subseteq V(G)$, that is, $V(H)=S$ and $E(H)=\{\{x,y\}\in E(G)\mid x,y\in S\}$. We say that $G$ is $H$-free if $H$ is not an induced subgraph of $G$. A hereditary class of graphs (or simply a class of graphs) is a family of graphs that is closed under taking induced subgraph.
Connectivity and block structures
A graph $G$ is connected if for every $x,y\in V(G)$, there exists a path from $x$ to $y$ in $G$. The connected components of $G$ are the maximal connected subgraphs of $G$. If $G$ has at least 2 vertices, a cut vertex of $G$ is a vertex $v$ such that $G\setminus\{v\}$ has strictly more connected components that $G$ does. For a graph on 1 vertex, in this paper, we define its only vertex to be a cut vertex.
Let $G$ be a graph and $r$ a cut vertex of $G$. Let $C_{1},\ldots,C_{k}$ be the connected components of $G\setminus\{r\}$ (notice that $k\geq 2$). For $1\leq i\leq k$, we set $G_{i}=G[V(C_{i})\cup\{r\}]$, that is, every $G_{i}$ is the graph obtained by adding back $r$ to $C_{i}$. Let $G^{\prime}=\bigoplus_{i=1}^{k}G_{i}$. Notice that for $1\leq i\leq k$, $G_{i}$ is exactly one connected component of $G^{\prime}$ and it contains a copy of $r$, that we denote by $r_{i}$. We call $G^{\prime}$ the split of $G$ over $r$ and we denote it by $\text{split}(G,r)$, we refer to $\{r_{1},\dots,r_{k}\}$ as the copies of $r$ in $G^{\prime}$.
We recall that every graph in this paper is finite, loopless, and without multiple edges. In this setting, a graph $G$ is 2-connected if for every $v\in V(G)$, the graph $G\setminus\{v\}$ is connected. A maximal 2-connected subgraph of $G$ is called a block of $G$. We call a vertex $v$ in a block $B$ of $G$ an internal vertex if it is not contained in any block of $G$ but $B$. We use the following properties of blocks of $G$ several times without necessarily referring to the following proposition. For the proof, see Proposition 5.3 in [BM08] and the paragraph under its proof.
Proposition 2.1.
For every graph $G$:
  1. (1)
    every two blocks of $G$ are either disjoint or have exactly one vertex in common,
  2. (2)
    the blocks of $G$ partition the edge-set of $G$,
  3. (3)
    every cycle of $G$ is contained in a block of $G$,
  4. (4)
    a vertex is an internal vertex of a block of $G$ if and only if it is not a cut vertex of $G$.
The blocks of a graph form a tree-like structure. Let us make this statement precise by discussing two graphs that capture the block structure of a graph $G$, namely the block tree and the block graph of $G$.
Let $G$ be a graph. We denote the set of its blocks by $\mathscr{B}(G)$ and the set of its cut vertices by $\mathscr{C}(G)$.
First, we define a graph $\mathcal{T}(G)$ (or $\mathcal{T}$ if the underlying graph is clear from the context) by setting $V(\mathcal{T})=\mathscr{B}(G)\sqcup\mathscr{C}(G)$ and $E(T)=\{\{c,b\}\mid b\in\mathscr{B}(G),c\in\mathscr{C}(G),c\in b\}$. It follows directly from Proposition 2.1 that $\mathcal{T}(G)$ is a tree and it is called the block tree of $G$ (see Section 5.2 of [BM08] for more). Notice that $\mathcal{T}$ always has a natural bipartition (or equivalently 2-colouring) into vertices that are blocks of $G$ and vertices that are cut vertices of $G$. Thus we will sometimes refer to the colour of the former as white and of the latter as black.
Second, we define a graph $\mathcal{B}(G)$ (or $\mathcal{B}$ if the underlying graph is clear from the context) by setting $V(\mathcal{B})=\mathscr{B}(G)$ and $E(\mathcal{B})=\{\{b,b^{\prime}\}\mid b\neq b^{\prime},b\cap b^{\prime}\neq\varnothing\}$, that is, $\mathcal{B}(G)$ is the intersection graph of the blocks of $G$. It is called the block graph of $G$.
A graph $H$ is called a block graph if there exists a graph $G$ such that $H$ is isomorphic to the block graph of $G$. There are numerous equivalent characterisations for block graphs. We use the following in this paper. The proof can be found in [Har63], Theorem A.
Proposition 2.2.
A graph $G$ is a block graph if and only if every block of $G$ is a complete graph.
Let us now define a function that will be used in several proofs. As discussed in the proof above, a block $b$ of $\mathcal{B}(G)$ contains at least two vertices and all these vertices have a common point, which is a cut vertex of $G$. Thus one can define a function $\mu=\mu_{G}:\mathscr{B}(\mathcal{B}(G))\to\mathscr{C}(G)$, which maps each block $B$ of of $\mathcal{B}(G)$ to the common vertex of vertices in $B$ (remember that the vertices of $B$ are blocks of $G$). That is, $\mu(B)=v$ where $\{v\}=\cap_{b\in B}b$.
Lemma 2.3.
$\mu$ is a bijection with inverse $\nu:\mathscr{C}(G)\to\mathscr{B}(\mathcal{B}(G))$ where $\nu(x)$ is the block of $\mathcal{B}(G)$ containing all the blocks of $G$ containing $x$. Moreover, for $x\in\mathscr{C}(G)$, $b\in\mathscr{B}(G)$, and $B\in\mathscr{B}(\mathcal{B}(G))$, we have that:
  • if $b\in B$, then $\mu(B)\in b$.
  • if $x\in b$, then $b\in\nu(x)$,
Proof.
First, if $\mu(B)=\mu(B^{\prime})=v$, then every vertex in $B$ and every vertex in $B^{\prime}$ have the vertex $v$ in common. Thus the subgraph of $\mathcal{B}$ induced on $B\cup B^{\prime}$ is a complete graph which is only possible if $B=B^{\prime}$. So, $\mu$ is injective. Second, for every cut vertex $x\in\mathscr{C}(G)$, there exists at least two blocks $b$ and $b^{\prime}$ of $G$ containing $x$. Therefore, $bb^{\prime}\in E(\mathcal{B}(G))$, so they are in a common block $B$ of $\mathcal{B}(G)$. So, $x=\mu(B)$, hence $\mu$ is surjective, so that it has an inverse. It is straightforward to see that $\nu:\mathscr{C}(G)\to\mathscr{B}(\mathcal{B}(G))$ defined in the statement is the inverse of $\mu$. The rest of the statement follows directly from the definitions of the two functions. ∎
Center of graphs
Every graph $G$ is equipped with a distance $d=d_{G}:V(G)\times V(G)\to\leavevmode\nobreak\ [0,+\infty]$ called the graph distance where for $x,y\in V(G)$, the distance $d(v,u)$ is the length of a shortest path from $v$ to $u$ in $G$. Notice that $G$ is connected if and only if $d_{G}(\cdot,\cdot)<+\infty$. For every vertex $v\in V(G)$, the eccentricity of $v$, denoted by $e(v)$, is defined to be the its distance to the vertex furthest from it, that is $e(v)=\max_{u\in V(G)}d(v,u)$. We denote by $E(v)$ the set $\{u\in V(G)\mid d(v,u)=e(v)\}$ and call each element of $E(v)$ an eccentric vertex for $v$.
The center of $G$, denoted by $Z(G)$, is the set of vertices with minimum eccentricity, that is $Z(G)=\{v\in V(G)\mid\forall u\in V(G)\ e(u)\geq e(v)\}$.
We recall the following theorem.
Theorem 2.4.
Let $G$ be a connected graph. There exists a block $B$ of $G$ such that $Z(G)\subseteq B$. Moreover, the center is either a singleton containing a cut vertex or there exists a unique block containing it.
Proof.
For the proof of the fact that the center is contained in a block, see Theorem 2.2 in [BH90]. If $Z(G)$ contains an internal vertex $u$, then Item (4) of Proposition 2.1 implies that there exists a unique block $B$ (namely the unique block containing $u$) such that $Z(G)\subseteq B$. If $Z(G)$ has no internal vertices, then either it contains exactly one cut vertex, in which case there is nothing to prove, or it contains at least two distinct cut vertices $v$ and $v^{\prime}$. In this case, let $B$ and $B^{\prime}$ be blocks containing $Z(G)$, then $v$ and $v^{\prime}$ are both in $B\cap B^{\prime}$, which, by Item (1) of Proposition 2.1 is only possible if $B=B^{\prime}$ and this completes the proof. ∎
The number of walks in graphs
Let $x$, $y$, $z\in V(G)$ for some graph $G$ and fix $i\geq 0$. We define $w_{i}(x,z)$ to be the number of walks of length $i$ from $x$ to $z$ in $G$. Recall that $w_{i}(x,z)$ is equal to $[A^{i}]_{xz}$, where $A$ is the adjacency matrix of $G$. We also define $w_{i}(x,z;y)$ to be the number of walks of length $i$ from $x$ to $z$ going through $y$.
For $i\geq 0$ and $x$, $y\in V(G)$, let a simple walk from $x$ to $y$ be a walk passing through $y$ only at the endpoints (notice it is not a symmetric notion). We let $w_{i}^{s}(x,y)$ denote the number of simple walks from $x$ to $y$.
Lemma 2.5.
For every $i\geq 0$, we have:
$w_{i}(x,z;y)=\sum_{j+k+l=i}w_{j}^{s}(x,y)w_{k}(y,y)w_{l}^{s}(y,z).$
Proof.
Observe that every walk of length $i\geq 0$ from $x$ to $z$ going through $y$ decomposes uniquely as the composition of a simple walk from $x$ to $y$ of length $j$, a walk from $y$ to $y$ of length $k$, and a simple walk from $y$ to $z$ of length $l$, where $j+k+l=i$. Conversely, given such a triple, their concatenation does give a walk of lenghth $i$ from $x$ to $z$ passing through $y$. Hence there is a bijection between the walks of length $i$ from $x$ to $z$ passing through $y$, and the triples $(p,q,r)$ where $p$ is a simple walk from $x$ to $y$ of length $j$, $q$ a walk from $y$ to $y$ of length $k$, and $r$ a simple walk from $y$ to $z$ of length $l$, with $j+k+l=i$. This leads to the desired formula. ∎
The following lemma, which will be needed later on, is based on a claim used in the proof of Theorem 6 of [KPS19]. We give it here with a detailed proof for the sake of completeness.
Lemma 2.6.
Let $G$ and $H$ be two graphs and let $x$, $y$, $z\in V(G)$, and $a$, $b$, $c\in V(H)$. Assume that for every $i\geq 0$, we have:
  1. (1)
    $w_{i}(x,y)=w_{i}(a,b)$,
  2. (2)
    $w_{i}(y,y)=w_{i}(b,b)$,
  3. (3)
    $w_{i}(y,z)=w_{i}(b,c)$.
Then for every $i\geq 0$ we have $w_{i}(x,z;y)=w_{i}(a,c;b)$.
Proof.
First, let us show that for all $i\geq 0$ we have $w_{i}^{s}(x,y)=w_{i}^{s}(a,b)$. We prove it by induction on $i\geq d_{G}(x,y)$. Indeed, the result is trivially true for $i<d_{G}(x,y)$. Now take $i=d_{G}(x,y)$. In this case, since $w_{j}(a,b)=w_{j}(x,y)=0$ for any $j<i$, and since $w_{i}(a,b)=w_{i}(x,y)>0$, we have that $i=d_{H}(a,b)$ as well. Then any path of length $i$ from $x$ to $y$ (or $a$ to $b$) is simple, so $w_{i}^{s}(x,y)=w_{i}(x,y)=w_{i}(a,b)=w_{i}^{s}(a,b)$, as desired.
Now assume that $w_{j}^{s}(x,y)=w_{j}^{s}(a,b)$ for any $j\leq i$ for some $i\geq d_{G}(x,y)$, and let us show that $w_{i+1}^{s}(x,y)=w_{i+1}^{s}(a,b)$. We have $w_{i+1}(x,y)=\sum_{j+k=i+1}w_{j}^{s}(x,y)w_{k}(y,y)$, and similarly for $w_{i+1}(a,b)$. Using the induction hypothesis, we reach:
$w_{i+1}^{s}(x,y)$ $=w_{i+1}^{s}(x,y)w_{0}(y,y)=w_{i+1}(x,y)-\sum_{j=0}^{i}w_{j}^{s}(x,y)w_{i+1-j}(y,y)$
$=w_{i+1}(a,b)-\sum_{j=0}^{i}w_{j}^{s}(a,b)w_{i+1-j}(b,b)$
$=w_{i+1}^{s}(a,b),$
as desired. This shows that for all $i\geq 0$ we have $w_{i}^{s}(x,y)=w_{i}^{s}(a,b)$. By symmetry, we also obtain that $w_{i}^{s}(y,z)=w_{i}^{s}(b,c)$ for all $i\geq 0$.
Finally, let $i\geq 0$. By Lemma 2.5, we have:
$w_{i}(x,z;y)$ $=\sum_{j+k+l=i}w_{j}^{s}(x,y)w_{k}(y,y)w_{l}^{s}(y,z)$
$=\sum_{j+k+l=i}w_{j}^{s}(a,b)w_{k}(b,b)w_{l}^{s}(b,c)$
$=w_{i}(a,c;b).$
This concludes the proof. ∎
Magic unitaries and quantum isomorphism
We now introduce quantum graph isomorphisms, the main object of this work. The definition is based on the notion of magic unitary matrix, introduced by Banica in [Ban05], extending work by Bichon on quantum automorphism groups of graphs [Bic03].
Let $X$ be a unital $C^{*}$-algebra. A magic unitary with coefficients in $X$ is a matrix $U=(u_{ij})_{i\in I,j\in J}$ where $I$ and $J$ are finite sets and such that for all $i\in I$ and all $j\in J$:
  1. (1)
    $u_{ij}^{2}=u_{ij}$ and $u_{ij}^{*}=u_{ij}$,
  2. (2)
    $\sum_{k\in I}u_{kj}=1=\sum_{k\in J}u_{ik}$.
Notice that (2) implies that $|I|=|J|$. Also, it is well-known that projections that form a partition of unity in a $C^{*}$-algebra are orthogonal, hence the rows and the columns of a magic unitary are orthogonal. This implies that $U$ is a unitary in $\mathcal{M}_{I,J}(X)$.
A quantum isomorphism from a graph $G$ to a graph $H$ with the same number of vertices is a magic unitary $U$ indexed by $V(H)\times V(G)$ such that $U\operatorname{Adj}(G)=\operatorname{Adj}(H)U$.
Mančinska and Roberson [MR19] proved a beautiful equivalent definition.
Theorem 2.7 (Theorem 7.16 in [MR19]).
Two graphs $G$ and $H$ are quantum isomorphic if and only if for every planar graph $P$ the number of morphisms from $P$ to $G$ is equal to the number of morphisms from $P$ to $H$.
The following lemma is a reformulation of a lemma due to Fulton, see Section 3.2 of [Ful06].
Lemma 2.8.
Let $U$ be a quantum isomorphism from $G$ to $H$. Let $x$, $y\in V(G)$ and $a$, $b\in V(H)$. If $u_{ax}u_{by}\neq 0$, then for all $i\geq 0$ we have $w_{i}(x,y)=w_{i}(a,b)$. In particular, we have $d(x,y)=d(a,b)$.
Proof.
Let $A=\operatorname{Adj}(G)$ and $B=\operatorname{Adj}(H)$. Let $i\geq 0$. We have $UA^{i}=B^{i}U$, so $\sum_{z\in V(G)}u_{az}[A^{i}]_{zy}=[UA^{i}]_{ay}=[B^{i}U]_{ay}=\sum_{c\in V(H)}[B^{i}]_{ac}u_{cy}$. Multiplying on the left by $u_{ax}$ and on the right by $u_{by}$ and using the orthogonality of the rows and columns of $U$, we obtain $u_{ax}u_{by}[A^{i}]_{xy}=[B^{i}]_{ab}u_{ax}u_{by}$. Since $u_{ax}u_{by}\neq 0$, we have that $[A^{i}]_{xy}=[B^{i}]_{ab}$. Recalling that $w_{i}(x,y)=[A^{i}]_{xy}$ and that $w_{i}(a,b)=[B^{i}]_{ab}$, we obtain the desired result.
Let us prove the last assertion by contrapositive. Up to symmetry, let us assume that $d(a,b)>d(x,y)=k\in\mathbb{N}$. Hence, we have $w_{k}(x,y)>0=w_{k}(a,b)$, so by what precedes we have $u_{ax}u_{yb}=0$, as desired. This concludes the proof. ∎
3. Quantum isomorphism of partitioned graphs and 2-connectedness
A partitioned graph is a pair $(G,\mathcal{P})$ where $G$ is a graph and $\mathcal{P}$ is a partition of $V(G)$. We denote by $\sim_{\mathcal{P}}$ the equivalence relation naturally associated to $\mathcal{P}$ on $V(G)$.
Given two partitioned graphs $(G,\mathcal{P}_{G})$ and $(H,\mathcal{P}_{H})$, we say that $U$ is a quantum isomorphism of partitioned graphs from $(H,\mathcal{P}_{H})$ to $(G,\mathcal{P}_{G})$ if, for all $x$, $y\in V(G)$ and $a$, $b\in V(H)$, we have $u_{xa}u_{yb}=0$ if $x\sim_{\mathcal{P}_{G}}y$ and $a\not\sim_{\mathcal{P}_{H}}b$, or $x\not\sim_{\mathcal{P}_{G}}y$ and $a\sim_{\mathcal{P}_{H}}b$.
Let $U$ be a such a quantum isomorphism with coefficients in a unital $C^{*}$-algebra $X$. Let $C\in\mathcal{P}_{G}$ and $D\in\mathcal{P}_{H}$. For $x\in V(G)$ and $a\in V(H)$, set $p_{C}(a)=\sum_{z\in C}u_{za}$ and $q_{D}(x)=\sum_{c\in D}u_{xc}$.
Lemma 3.1.
The functions $p_{C}\colon V(H)\to X$ and $q_{D}\colon V(G)\to X$ are projection-valued and constant on the cells of $\mathcal{P}_{H}$ and $\mathcal{P}_{G}$ respectively.
Proof.
First, using the fact that rows and columns of $U$ are orthogonal, it is easy to see that $p_{C}$ and $q_{D}$ are projection-valued.
Let $a$, $b\in V(H)$ be in a common cell of $\mathcal{P}_{H}$. We have:
$p_{C}(a)(1-p_{C}(b))$ $=\left(\sum_{x\in C}u_{xa}\right)\left(\sum_{y\in V(H)}u_{yb}-\sum_{y\in C}u_{yb}\right)$
$=\sum_{x\in C,y\notin C}u_{xa}u_{yb}$
$=0$
since $U$ preserves the partitions. This shows that $p_{C}(a)p_{C}(b)=p_{C}(a)$. Taking adjoints, one obtains that $p_{C}(a)=p_{C}(b)p_{C}(a)$. Applying what precedes to $b$ and $a$, we reach $p_{C}(b)=p_{C}(b)p_{C}(a)=p_{C}(a)$, as desired.
Finally, notice that $U^{*}$ is a quantum isomorphism from $G$ to $H$ preserving the partitions. Adding the magic unitary used in index, we have that $q_{C,U}=p_{D,U^{*}}$, which is constant on cells of $\mathcal{P}_{G}$ by what precedes. This concludes the proof. ∎
We can now define $p_{CD}=\sum_{z\in C}u_{za}$ for some $a\in D$, and $q_{DC}=\sum_{c\in D}u_{xc}$ for some $x\in C$. When necessary, we can write $p_{CD}(U)$ and $q_{DC}(U)$ to recall the magic unitary used to build these projections. Recall that $q_{CD}(U^{*})=p_{CD}(U)$.
Lemma 3.2.
For any $C\in\mathcal{P}_{G}$ and $D\in\mathcal{P}_{H}$, we have $q_{DC}=p_{CD}$.
Proof.
Let $x\in C$ and $a\in D$. We have:
$q_{DC}(1-p_{CD})$ $=\left(\sum_{b\in D}u_{xb}\right)\left(\sum_{y\in V(G)}u_{ya}-\sum_{y\in C}u_{ya}\right)$
$=\sum_{b\in D,y\notin C}u_{xb}u_{ya}$
$=0$
since $a\sim_{\mathcal{P}_{H}}b$, $x\not\sim_{\mathcal{P}_{G}}y$, and $U$ preserves the partitions. Therefore, $q_{DC}=q_{DC}p_{CD}$. Applying what precedes to $q_{CD}(U^{*})$ and $p_{DC}(U^{*})$, we reach:
$p_{CD}(U)$ $=q_{CD}(U^{*})=q_{CD}(U^{*})p_{DC}(U^{*})$
$=p_{CD}(U)q_{DC}(U)=q_{DC}(U)^{*}$
$=q_{DC}(U),$
as desired. This concludes the proof. ∎
If $C\subseteq V(H)$ and $D\subseteq V(G)$, we denote by $U[C,D]$ the submatrix of $U$ indexed by $C\times D$. We now reach the fundamental theorem for partitioned graphs.
Theorem 3.3.
Let $(G,\mathcal{P}_{G})$ and $(H,\mathcal{P}_{H})$ be two partitioned graphs and denote by $A_{G}$ and $A_{H}$ respectively the adjacency matrices of $G$ and $H$. Let $U$ be a quantum isomorphism of partitioned graphs from $(G,\mathcal{P}_{G})$ to $(H,\mathcal{P}_{H})$ with coefficients in a unital $C^{*}$-algebra $X$. Take $C\in\mathcal{P}_{H}$ and $D\in\mathcal{P}_{G}$, let $W=U[C,D]$, and $p=p_{CD}$. Then:
  1. (1)
    the matrix $P=(p_{KL})_{K\in\mathcal{P}_{H},L\in\mathcal{P}_{G}}$ is a magic unitary with coefficients in $X$,
  2. (2)
    if $p\neq 0$, then it is the unit of the $C^{*}$-subalgebra of $X$ generated by the coefficients of $W$,
  3. (3)
    if $p=0$, then $W=0$,
  4. (4)
    $W$ is a magic unitary with coefficients in the $C^{*}$-algebra $pXp$,
  5. (5)
    $A_{H}[C,C]W=WA_{G}[D,D]$.
In particular, there is a bijection $\varphi\colon\mathcal{P}_{G}\to\mathcal{P}_{H}$ such that $W$ is a quantum isomorphism from $G[D]$ to $H[\varphi(D)]$ for every $D\in\mathcal{P}_{G}$.
Proof.
Let us start with (1). By Lemma 3.1, we have that $P$ is a well-defined matrix of projections. Let us check that $P$ sums to $1$ on columns. For $D\in\mathcal{P}_{G}$, taking $a\in D$, we have:
$\sum_{C\in\mathcal{P}_{H}}p_{CD}=\sum_{C\in\mathcal{P}_{H}}\sum_{x\in C}u_{xa}=\sum_{x\in V(G)}u_{xa}=1.$
Now for $C\in\mathcal{P}_{H}$, using Lemma 3.2, we have $p_{CD}(U)=q_{DC}(U)=p_{DC}(U^{*})$, so
$\sum_{K\in\mathcal{P}_{G}}p_{CK}(U)=\sum_{K\in\mathcal{P}_{G}}p_{KC}(U^{*})=1$
by the previous computation. Hence $P$ is a magic unitary, which proves (1).
(2) and (3) follow from the orthogonality of rows and columns of magic unitaries, from the construction of $p_{CD}$, and from Lemma 3.2.
(4) is a rewriting of the previous properties.
Let us check (5). By assumption, we have $UA_{G}=A_{H}U$, so $\sum_{K\in\mathcal{P}_{G}}U[C,K]A_{G}[K,D]=\sum_{K\in\mathcal{P}_{H}}A_{H}[C,K]U[K,D]$. Let $p=p_{CD}$. Since $P$ has orthogonal rows and columns, we have that $pU[C,K]=0$ when $K\neq D$, and $U[K,D]p=0$ when $K\neq C$. Since $A_{G}$ and $A_{H}$ are scalar-valued, multiplying the previous equation on the left and the right by $p$, we obtain $(pU[C,D]p)A_{G}[D,D]=A_{H}[C,C](pU[C,D]p)$. Finally, notice that $pU[C,D]p=U[C,D]$ by (2) and (3). Hence we have the desired result.
It remains to prove the last statement. For this, applying Lemma 2.20 from [Meu23] to the magic unitary $P$ we have a bijection $\varphi\colon\mathcal{P}_{G}\to\mathcal{P}_{H}$ such that $p_{\varphi(D)D}\neq 0$ for every $D\in\mathcal{P}_{G}$. Let $D\in\mathcal{P}_{G}$ and set $C=\varphi(D)$ and $W=U[C,D]$. By (4), $W$ is a magic unitary, and by (5) we have $A_{H}[C,C]W=WA_{G}[D,D]$. But $A_{G}[D,D]=\operatorname{Adj}(G[D])$ and $A_{H}[C,C]=\operatorname{Adj}(H[C])$. Hence $W$ is a quantum isomorphism from $G[D]$ to $H[C]$, as desired. This concludes the proof. ∎
The partition into connected components is automatically preserved.
Lemma 3.4.
Let $U$ be a quantum isomorphism from $G$ to $H$. Let $\mathcal{P}_{G}$ be the partition of $V(G)$ into the vertex sets of the connected components of $G$ and define $\mathcal{P}_{H}$ similarly. Then $U$ is a quantum isomorphism of partitioned graphs from $(G,\mathcal{P}_{G})$ to $(H,\mathcal{P}_{H})$. As a consequence, the connected components of $G$ and $H$ are in bijection and are two-by-two quantum isomorphic.
Proof.
The last assertion follows immediately from Theorem 3.3 once $U$ is shown to be a quantum isomorphism of partitioned graphs. For this, let $x$, $y\in V(G)$ and let $a$, $b\in V(H)$. Assume that $x$ and $y$ are in a common connected component of $G$ and that $a$ and $b$ are not in a common connected component of $H$. Then $d_{G}(x,y)<+\infty=d_{H}(a,b)$, so by Lemma 2.8 we have that $u_{ax}u_{by}=0$. Similarly, $u_{ax}u_{by}=0$ if $d_{G}(x,y)=+\infty>d_{H}(a,b)$. This concludes the proof. ∎
The following lemma relates the coefficients of a quantum isomorphism from $H$ to $G$ to the blocks of $G$ and $H$.
Lemma 3.5.
Let $U$ be a quantum isomorphism from $H$ to $G$. Let $x,y\in V(G)$ and $a,b\in V(H)$. If $x$ and $y$ are in a common block but not $a$ and $b$, or if $a$ and $b$ are in a common block but not $x$ and $y$, then $u_{xa}u_{yb}=0.$
Moreover, if $x$ is a cut vertex and $a$ is not, or $a$ is a cut vertex and $x$ is not, then $u_{xa}=0=u_{ax}$.
Proof.
Let $a$, $b\in V(H)$ not be in a common block, and take $x$, $y\in V(G)$ such that $u_{xa}u_{yb}\neq 0$. We want to show that $x$ and $y$ are not in a common block neither.
By assumption, there is a cut vertex $c\in V(H)$ such that every walk from $a$ to $b$ passes through $c$. In other words, we have $w_{i}(a,b;c)=w_{i}(a,b)$ for all $i\geq 0$. Now, we have
$0\neq u_{xa}1u_{yb}=u_{xa}\left(\sum_{z\in V(G)}u_{zc}\right)u_{yb}=\sum_{z\in V(G)}u_{xa}u_{zc}u_{yb},$
hence there exists $z\in V(G)$ such that $u_{xa}u_{zc}u_{yb}\neq 0$. In particular, we have $u_{xa}u_{zc}\neq 0$, $u_{zc}\neq 0$, and $u_{zc}u_{yb}\neq 0$, so by Lemma 2.8 for all $i\geq 0$ we have $w_{i}(x,z)=w_{i}(a,c)$, $w_{i}(z,z)=w_{i}(c,c)$, and $w_{i}(z,y)=w_{i}(c,b)$. Hence, by Lemma 2.6, we have $w_{i}(x,y;z)=w_{i}(a,b;c)=w_{i}(a,b)$ since $c$ is a cut vertex separating $a$ and $b$. Since by assumption $u_{xa}u_{yb}\neq 0$, by Lemma 2.8 again we have that $w_{i}(x,y)=w_{i}(a,b)$, so we have shown that $w_{i}(x,y)=w_{i}(x,y;z)$. This implies that $z$ is a cut vertex separating $x$ and $y$, hence $x$ and $y$ are not in the same block of $G$.
Applying what precedes to $U^{*}$ and $H$ and $G$, we obtain that if $u_{xa}u_{yb}\neq 0$ and $x$ and $y$ are not in a common block, then neither are $a$ and $b$. This concludes the proof of the first part of the lemma.
Now let $x$ be a cut vertex in $G$ and take $a\in V(H)$ such that $u_{xa}\neq 0$. By definition, there are $y$ and $z\in V(G)$ which are neighbors of $x$ but are not in a common block of $G$. Now, we have:
00 0 $\neq u_{xa}=u_{xa}^{2}$
$=u_{xa}\left(\sum_{b\in V(H)}u_{yb}\right)\left(\sum_{c\in V(H)}u_{zc}\right)u_{xa}$
$=\sum_{b,c\in V(H)}u_{xa}u_{yb}u_{zc}u_{xa}.$
In particular, there exist $b$, $c\in V(H)$ such that $u_{xa}u_{yb}u_{zc}u_{xa}\neq 0$. This implies that $u_{xa}u_{yb}\neq 0$ and that $u_{zc}u_{xa}\neq 0$, so by Lemma 2.8 we have that $ab\in E(H)$ and $ca\in E(H)$. Moreover, we have that $u_{yb}u_{zc}\neq 0$, so by what precedes $b$ and $c$ are not in a common block of $H$, since $y$ and $z$ are not. Hence, $b$ and $c$ are neighbors of $a$ which are not in a common block of $H$. So $a$ is in the intersection of two distinct blocks, thus it is a cut vertex. Reasoning similarly with $U^{*}$, we obtain that if $a$ is a cut vertex and $u_{xa}\neq 0$, then $x$ is a cut vertex. This concludes the proof. ∎
Theorem 3.6.
If $G$ is 2-connected and quantum isomorphic to $H$, then $H$ is 2-connected.
Proof.
Let us prove it by contrapositive. Assume that $H$ is not 2-connected. Hence there is a cut vertex $a\in V(H)$. Let $U$ be a quantum isomorphism from $G$ to $H$. Since $1=\sum_{x\in V(G)}u_{ax}$, there is $x\in V(G)$ such that $u_{ax}\neq 0$. Hence $x$ is a cut vertex by Lemma 3.5, so $G$ is not 2-connected. This concludes the proof. ∎
Lemma 3.7.
Let $G$ and $H$ be quantum isomorphic graphs and let $U$ be a quantum isomorphism from $H$ to $G$. Then:
  • $Z(G)$ is a cut vertex of $G$ if and only if $Z(H)$ is a cut vertex of $H$, in which case $u_{z_{G}z_{H}}=1$, where $\{z_{G}\}=Z(G)$ and $\{z_{H}\}=Z(H)$,
  • if $Z(G)$ is a not a cut vertex, then neither is $Z(H)$, and by Theorem 2.4 we can let $B_{Z}(G)$ be the unique block of $G$ containing $Z(G)$, we define similarly $B_{Z}(H)$. Then $U[B_{Z}(G),B_{Z}(H)]$ is a quantum isomorphism from $B_{Z}(H)$ to $B_{Z}(G)$.
Proof.
Assume that $Z(G)=\{z_{G}\}$ is a cut vertex. By Theorem 3.4 of [Meu23], $U[Z(G)],Z(H)]$ is a quantum isomorphism from $Z(H)$ to $Z(G)$, hence $Z(H)=\{z_{H}\}$ is a singleton, and $u_{z_{G}z_{H}}=1$. In particular, by Lemma 3.5, $z_{H}$ is a cut vertex in $H$, as desired. We obtain the reverse direction by exchanging $G$ and $H$ and applying what precedes to $U^{*}$.
Now assume that $Z(G)$ is not a cut vertex. By what precedes, neither is $Z(H)$. Let $a\in B_{Z}(H)$ and let $x\in V(G)$ such that $u_{xa}\neq 0$. Let $b\in Z(H)$. We have $0\neq u_{xa}=u_{xa}\left(\sum_{y\in V(G)}u_{yb}\right)=\sum_{y\in V(G)}u_{xa}u_{yb}$, so there exists $y\in V(G)$ such that $u_{xa}u_{yb}\neq 0$. In particular, $u_{yb}\neq 0$ and $b\in Z(H)$, so by Corollary 3.3 of [Meu23] $y\in Z(G)$. By definition, $a$ and $b$ are in a common block of $G$, so by Lemma 3.5 $x$ and $y$ are in a common block of $G$ as well. Since $y\in Z(G)$, we have that $x\in B_{Z}(G)$, as desired.
Hence we have shown that if $a\in B_{Z}(H)$ and $x\notin B_{Z}(G)$, then $u_{xa}=0$. Exchanging $G$ and $H$ and applying what precedes to $U^{*}$ leads to $u_{xa}=0$ when $x\in B_{Z}(G)$ and $a\notin B_{Z}(H)$. Hence $U$ is diagonal by block, with $U=\operatorname{Diag}(U_{1},U_{2})$, where $U_{2}=U[B_{Z}(G),B_{Z}(H)]$ and $U_{1}=U[V(G)\setminus B_{Z}(G),V(H)\setminus B_{Z}(H)]$. This implies that $U_{2}$ is a magic unitary, and, since $U$ is a quantum isomorphism from $H$ to $G$, we have that $U_{2}$ is one from $B_{Z}(H)$ to $B_{Z}(G)$. This concludes the proof. ∎
4. Operations on anchored graphs
As explained in the introduction, our strategy is to work inductively by reducing to graphs whose connected components have fewer blocks that the original graph. This requires a formalism that we call anchored graphs which we will now introduce. We will then define and study the reduction operation, called the $\Gamma$ operation.
4.1. Anchored graphs
Definition 4.1.
A connected anchored graph is a pair $(G,R)$ where $G$ is a connected graph and $R\subseteq V(G)$ is either a cut vertex of $G$ or a block of $G$. An anchored graph is a pair $(G,R)$ where $G$ is a graph, $R\subseteq V(G)$, and for every connected component $G_{i}$ of $G$, the pair $(G_{i},R\cap V(G_{i}))$ is a connected anchored graph.
Let $(G,R)$ and $(H,S)$ be anchored graphs. An isomorphism from $(G,R)$ to $(H,S)$ is a graph isomorphism $\phi:V(G)\to V(H)$ from $G$ to $H$ that preserves the anchors, that is $\phi(R)=S$. A quantum isomorphism from $(G,R)$ to $(H,S)$ is a magic unitary $U$ such that $U\operatorname{Adj}(G)=\operatorname{Adj}(H)U$ and that preserves the anchors, that is $u_{ax}=0$ if $x\in R$ and $a\notin S$ or if $x\notin R$ and $a\in S$.
The following lemma can be seen as an anchored version of Theorem 3.7 of [Meu23].
Lemma 4.2.
Let $(G,R)$ and $(H,S)$ be two quantum isomorphic anchored graphs. Let $G_{1},\dots,G_{k}$ be the connected components of $G$ and $H_{1},\dots,H_{l}$ be the connected components of $H$. Let $R_{i}=R\cap V(G_{i})$ for $1\leq i\leq k$ and $S_{j}=S\cap V(H_{j})$ for $1\leq j\leq l$. Then $l=k$ and up to reordering $(G_{i},R_{i})$ is quantum isomorphic to $(H_{i},S_{i})$ for every $1\leq i\leq k$.
Proof.
Let $U$ be a quantum isomorphism of anchored graphs from $(G,R)$ to $(H,S)$. Let $\mathcal{P}_{G}$ be the partition of $V(G)$ into the sets of vertices of the connected components of $G$, and let $\mathcal{P}_{H}$ be similarly defined for $H$. By Lemma 3.4, $U$ is a quantum isomorphism of partitioned graphs from $(G,\mathcal{P}_{G})$ to $(H,\mathcal{P}_{H})$, so, by Theorem 3.3, we have that $k=l$ and that up to reordering $W_{i}=U[V(H_{i}),V(G_{i})]$ is a quantum isomorphism from $G_{i}$ to $H_{i}$. We claim it is a quantum isomorphism of anchored graphs from $(G_{i},R_{i})$ to $(H_{i},S_{i})$. Indeed, let $x\in V(G_{i})$ and $a\in V(H_{i})$ such that $w_{xa}\neq 0$. If $x\in R_{i}$, then $x\in R$ so since $U$ is a quantum isomorphism of anchored graphs and $u_{xa}=w_{xa}\neq 0$ we have $a\in R$. Since $a\in V(H_{i})$, we have $a\in S_{i}$, as desired. Similarly, we obtain that if $a\in S_{i}$, then $x\in R_{i}$. Hence $W_{i}$ is a quantum isomorphism of anchored graphs from $(G_{i},R_{i})$ to $(H_{i},S_{i})$ for every $1\leq i\leq k$, which concludes the proof. ∎
Let $G$ be a graph. Define a function $\rho=\rho_{G}:V(G)\to\mathscr{P}(V(G))$ that maps $v$ to $\{v\}$ if $v$ is a cut vertex or an isolated vertex of $G$ and to the vertex set of the unique block containing $v$ otherwise. Notice that we always have $v\in\rho(v)$.
4.2. Block structures of anchored graphs
A connected rooted graph is a pair $(F,r)$, where $F$ is a graph and $r\in V(F)$. Since all the graphs that we will consider as rooted graphs are connected, we refer to them simply as rooted graphs. An isomorphism $\alpha$ from a rooted graph $(F,r)$ to a rooted graph $(F^{\prime},r^{\prime})$ is an isomorphism from $F$ to $F^{\prime}$ such that $\alpha(r)=r^{\prime}$.
Let $(G,R)$ be a connected anchored graph. Define its block tree to be the rooted tree $\mathcal{T}(G,R)=(\mathcal{T}(G),R)$. Notice that $R$ is always a single vertex of $\mathcal{T}(G)$ and that in case that $R=\{r\}$ is a cut vertex of $G$, the correct notation would be $(\mathcal{T}(G),r)$, however, by abuse of notation we keep writing $(\mathcal{T}(G),R)$.
Also, define its block graph to be the anchored graph $\mathcal{B}(G,R)=(\mathcal{B}(G),\tilde{R})$ where $\tilde{R}=\rho_{\mathcal{T}(G)}(R)$ if $R$ is not a cut vertex of $G$ and $\tilde{R}=\nu_{G}(R)$ if $R$ is a cut vertex. In other words, the underlying graph of $\mathcal{B}(G,R)$ is the block graph of $G$ and it is turned into an anchored graph by considering the block corresponding to the cut vertex in $R$, when $R$ consists of a single cut vertex of $G$, or the block containing the vertex $R\in V(\mathcal{B}(G))$ when $R$ is a block of $G$, or $R$ itself when it is already a cut vertex in $\mathcal{B}(G)$.
Lemma 4.3.
Assume that the rooted trees $\mathcal{T}(G,R)$ and $\mathcal{T}(H,S)$ are isomorphic. Then $R\in\mathscr{B}(G)$ if and only if $S\in\mathscr{B}(H).$
Proof.
Let $(K,L)$ be an anchored graph and let $\mathcal{T}=\mathcal{T}(K,L)$. Notice that by construction, $\mathcal{T}$ is a bipartite rooted tree whose leaves are white. Recall that the depth of $\mathcal{T}$ is the quantity $d(\mathcal{T})=\max_{s\in V(\mathcal{T})}d(L,s)$. Hence, an immediate induction shows that the colour of the root $L$ is white if $d(\mathcal{T})$ is odd, and black if $d(\mathcal{T})$ is even. Since the depth of a rooted tree is preserved under isomorphism, applying what precedes to $\mathcal{T}(G,R)$ and $\mathcal{T}(H,S)$, we obtain that $R$ is white if and only if $S$ is white, that is, $R\in\mathscr{B}(G)$ if and only if $S\in\mathscr{B}(H)$, as desired. ∎
4.3. Operation $\Gamma$
Given a connected anchored graph $(G,R)$, we want to define $\Gamma(G,R)=(G^{\prime},R^{\prime})$, another anchored graph, where the number of blocks of each connected component of $G^{\prime}$ is strictly less than that of $G$. This operation will be key to the inductive processes in the proofs of our main results. We start with a formal definition.
Definition 4.4.
Let $(G,R)$ be a connected anchored graph. We define
$\Gamma(G,R)=\begin{cases}(\text{split}(G,r),\rho_{\text{split}(G,r)}(\{r_{1},\dots,r_{k}\}))&\text{ if }R=\{r\}\text{ is a cut vertex of }G,\\ (G\setminus E(G[R]),\rho_{G\setminus E(G[R]}(R))&\text{ if }R\text{ is not a cut vertex of }G.\end{cases}$
Let us make this definition more explicit:
Case 1: $R=\{r\}$ where $r$ is a cut vertex of $G$. Let $C_{1},\ldots,C_{k}$ be the connected components of $G\setminus\{r\}$ (notice that $k\geq 2$). For $1\leq i\leq k$, we set $G_{i}=G[V(C_{i})\cup\{r\}]$, that is, every $G_{i}$ is the graph obtained by adding back $r$ to $C_{i}$. Let $G^{\prime}=\bigoplus_{i=1}^{k}G_{i}$. Notice that for $1\leq i\leq k$, $G_{i}$ is exactly one connected component of $G^{\prime}$ and it contains a copy of $r$, that we denote by $r_{i}$. Set $R^{\prime}=\rho(\{r_{1},\dots,r_{k}\}$). Notice that by construction, no $r_{i}$ is a cut vertex in $G_{i}$. So, $R_{i}=\rho(r_{i})$ is the vertex set of the unique block of $G_{i}$ containing $r_{i}$, and $R^{\prime}=\cup_{i=1}^{k}R_{i}$. Set $\Gamma(G)=(G^{\prime},R^{\prime})$.
Case 2: $R$ is a block of $G$. Define $G^{\prime}$ to be the graph obtained from $G$ by removing all edges in $R$. That is, $G^{\prime}=G\setminus\{xy\in E(G)\mid x,y\in R\}$. We also set $R^{\prime}=\rho(R)$. Set $\Gamma(G)=(G^{\prime},R^{\prime})$.
Notice that by construction if $(G^{\prime},R^{\prime})=\Gamma(G,R)$, we have that every connected component of $G^{\prime}$ has strictly less blocks than $G$.
Lemma 4.5.
Let $(G,R)$ be a connected anchored graph and assume that $R$ is not a cut vertex of $G$. Set $(G^{\prime},R^{\prime})=\Gamma(G,R)$. Notice that $V(G)=V(G^{\prime})$. We have:
  1. (1)
    $R\subseteq R^{\prime}$,
  2. (2)
    if $v\in R^{\prime}\setminus R$, then there exists $u\in R$ such that $v$ and $u$ are in a common block of $G$, and the number of blocks of $G$ that contain $u$ is exactly 2,
  3. (3)
    if $u\in R$ and the number of blocks of $G$ that contain $u$ is exactly 2, then for every $v\in V(G)$ such that $u$ and $v$ are in a common block of $G$, we have $v\in R^{\prime}$.
Proof.
Let $\rho=\rho_{G^{\prime}}$. The first item follows directly from the fact that for every $v$, we have $v\in\rho(v)$.
If $v\in R^{\prime}$, then there exists $u\in R$ such that $v\in\rho(u)$. Now, if $v\in R^{\prime}\setminus R$, then $v\neq u$, and in particular we cannot have $\rho(u)=\{u\}$. That is, $\rho(u)$ is the vertex set of the unique block $B$ of $G^{\prime}$ containing $u$. Thus $v\in B$ as well. Moreover, $\rho(u)\neq\{u\}$ implies that $u$ is neither a cut vertex nor an isolated vertex of $G^{\prime}$. Firstly, since it is not a cut vertex of $G$, it is in at most 2 blocks of $G$. Secondly, since it is not an isolated vertex of $G^{\prime}$, it is in at least two blocks of $G$. This completes the proof of (2).
Finally, let $u$ be a vertex as described in the statement of (3). Let $B$ be the block of $G$ other than $G[R]$ containing $u$. By removing the edges of $G[R]$, this vertex $u$ will be contained in exactly one block $B$ of $G$. Thus $\rho(u)=V(B)$. Now, let $v$ be a vertex in a common block with $u$. If $v\in R$, then the result follows from (1). If $v\in B$, then $v\in\rho(u)\subseteq R^{\prime}$ which completes the proof. ∎
Now given a quantum isomorphism $U$ from $(G,R)$ to $(H,S)$, we would like to find a quantum isomorphism $\Gamma(U)$ from $\Gamma(G,R)$ to $\Gamma(H,S)$. We start by checking that in this case $R$ and $S$ are of the same nature.
Lemma 4.6.
Let $U$ be a quantum isomorphism from $(G,R)$ to $(H,S)$. Then $U[S,R]$ is a quantum isomorphism from $G[R]$ to $H[S]$. Moreover, if $R$ is a cut vertex, so is $S$, and if $R$ is a block, so is $S$.
Proof.
By definition, $U=\operatorname{Diag}(U_{1},U_{2})$ is diagonal by block with $U_{2}=U[S,R]$. In particular, $U[S,R]$ is a magic unitary, and since $U\operatorname{Adj}(G)=\operatorname{Adj}(H)U$, we have $U_{2}\operatorname{Adj}(G[R])=U_{2}\operatorname{Adj}(G)[R,R]=\operatorname{Adj}(H)[S,S]U_{2}=\operatorname{Adj}(H[S])U_{2}$, so $U_{2}$ is a quantum isomorphism from $G[R]$ to $H[S]$. The fact that $S$ is a cut vertex if and only if $R$ is then follows from Lemma 3.5. This concludes the proof. ∎
When $R$ is not a cut vertex of $G$, we set $\Gamma(U)=U$. When $R=\{r\}$ is a cut vertex of $G$, we have $U=\operatorname{Diag}[U_{0},1]$. We define $\Gamma(U)=V=\operatorname{Diag}[U_{0},P(U_{0})]$. That is, after choosing $w_{j}\in C_{j}$ for every $j$, we set:
$v_{xy}=\begin{cases}u_{xy}&\text{ if }x\notin\{s_{1},\dots,s_{k}\}\text{ and }y\notin\{r_{1},\dots,r_{k}\}\\ \sum_{z\in D_{i}}u_{zw_{j}}&\text{ if }x=s_{i},y=r_{j}\\ 0&\text{ otherwise}\end{cases}.$
We will prove that $V$ is indeed a well-defined magic unitary.
Lemma 4.7.
Let $U$ be a quantum isomorphism from $(G,R)$ to $(H,S)$. Then $\Gamma(U)$ is a well-defined magic unitary.
Proof.
If $R$ is a block, then so is $S$ by Lemma 4.6. In this case, $\Gamma(U)=U$ and the result is clear. Let us assume that $R=\{r\}$ and $S=\{s\}$, with both $r$ and $s$ cut-vertices. Let $H_{0}=H\setminus\{s\}$ and $G_{0}=G\setminus\{r\}$. By definition, we have $U=\operatorname{Diag}(U_{0},1)$, where $U_{0}=U[V(H_{0}),V(G_{0})]$ and $u_{sr}=1$. Hence $U_{0}$ is a magic unitary, and is actually a quantum isomorphism from $G_{0}$ to $H_{0}$.
Notice that $G_{0}=\bigoplus_{i=1}^{k}C_{i}$ and $H_{0}=\bigoplus_{j=1}^{l}D_{j}$ for some $k$, $l\geq 1$ (actually we have $k=l\geq 2$ but we will not need it here). $\mathcal{C}=\{V(C_{1}),\ldots,V(C_{k})\}$ and $\mathcal{D}=\{V(D_{1}),\ldots,V(D_{l})\}$ are partitions of $V(G)$ and $V(H)$. By Lemma 3.4, we know that $U_{0}$ is a quantum isomorphism of partitioned graphs from $(G_{0},\mathcal{C})$ to $(H_{0},\mathcal{D})$. Hence by Theorem 3.3 we have that $P(U_{0})$ is a well-defined magic unitary. This implies that $\Gamma(U)=\operatorname{Diag}(U_{0},P(U_{0}))$ is a well-defined magic unitary. ∎
Lemma 4.8.
If $U$ is a quantum isomorphism from $(G,R)$ to $(H,S)$, then $\Gamma(U)$ is a quantum isomorphism from $\Gamma(G,R)$ to $\Gamma(H,S)$.
Proof.
Set $(G^{\prime},R^{\prime})=\Gamma(G,R)$ and $(H^{\prime},S^{\prime})=\Gamma(H,S)$.
Let us start with the case where $R=\{r\}$ and $r$ is a cut vertex in $G$. In this case, by Lemma 4.6, we have that $S=\{s\}$, with $s$ a cut vertex in $H$. Denote the copies of $r$ in $G^{\prime}=\text{split}(G,r)$ by $\{r_{1},\dots,r_{k}\}$ and the copies of $s$ in $H^{\prime}=\text{split}(H,s)$ by $\{s_{1},\dots,s_{l}\}$. Let $A$ and $B$ denote the adjacency matrices of $G$ and $H$ respectively. We have
$A=\begin{pmatrix}A_{1}&0&\ldots&0&&{}^{t}L_{1}\\ 0&A_{2}&\ddots&\vdots&&{}^{t}L_{2}\\ \vdots&\ddots&\ddots&0&&\vdots\\ 0&\ldots&0&A_{k}&&{}^{t}L_{k}\\ &&&&&\\ L_{1}&L_{2}&\ldots&L_{k}&&0\end{pmatrix},$
where $A_{j}$ is the adjacency matrix of $C_{j}$ (the $j$th connected component of $G\setminus R$) and the last line of $A$ is indexed by $r$. Setting $A_{0}=\operatorname{Diag}[A_{1},\dots,A_{k}]\text{ and }L=\operatorname{Diag}[L_{1},\dots,L_{k}],$ we have that the adjacency matrix of $G^{\prime}$ is
$A^{\prime}=\operatorname{Adj}(G^{\prime})=\begin{pmatrix}A_{0}&{}^{t}L\\ L&0\end{pmatrix}.$
Similarly,
$B=\begin{pmatrix}B_{1}&0&\ldots&0&&{}^{t}K_{1}\\ 0&B_{2}&\ddots&\vdots&&{}^{t}K_{2}\\ \vdots&\ddots&\ddots&0&&\vdots\\ 0&\ldots&0&B_{l}&&{}^{t}K_{l}\\ &&&&&\\ K_{1}&K_{2}&\ldots&K_{l}&&0\end{pmatrix},$
where $B_{i}$ is the adjacency matrix of $D_{i}$ (the $i$th connected component of $H\setminus S$) and the last line of $B$ is indexed by $s$. Again, setting $B_{0}=\operatorname{Diag}[B_{1},\dots,B_{l}]\text{ and }K=\operatorname{Diag}[K_{1},\dots,K_{l}],$ we get that the adjacency matrix of $H^{\prime}$ is
$B^{\prime}=\operatorname{Adj}(H^{\prime})=\begin{pmatrix}B_{0}&{}^{t}K\\ K&0\end{pmatrix}.$
Using the notations from the definition of $\Gamma(U)$, we have $V=\Gamma(U)=\operatorname{Diag}[U_{0},P(U_{0})]$, where
$U_{0}=\begin{pmatrix}U_{11}&\ldots&U_{1k}\\ \vdots&\ddots&\vdots\\ U_{l1}&\ldots&U_{lk}\end{pmatrix}.$
In particular, since $U_{0}$ is a magic unitary, we have $k=l$. Here, $U_{ij}$ is the submatrix of $U$ indexed by rows corresponding to $D_{i}$ and columns corresponding to $C_{j}$, that is $U_{ij}=U[D_{i},C_{j}]$. Let $P=P(U_{0})$, it is a $k\times k$ matrix whose entry in row $i$ and column $j$ is defined to be $p_{ij}=\sum_{z\in D_{i}}u_{zw_{i}}$. Remember that by Theorem 3.3 it is equal to the sum of any row or any column of $U_{ij}$. We have
$VA^{\prime}=\begin{pmatrix}U_{0}&0\\ 0&P\end{pmatrix}\begin{pmatrix}A_{0}&{}^{t}L\\ L&0\end{pmatrix}=\begin{pmatrix}U_{0}A_{0}&U_{0}{}^{t}L\\ PL&0\end{pmatrix},$
and
$B^{\prime}V=\begin{pmatrix}B_{0}&{}^{t}K\\ K&0\end{pmatrix}\begin{pmatrix}U_{0}&0\\ 0&P\end{pmatrix}=\begin{pmatrix}B_{0}U_{0}&{}^{t}KP\\ KU_{0}&0\end{pmatrix}.$
Since $UA=BU$, we have $A_{0}U_{0}=U_{0}B_{0}.$ Now, let us verify that $KU_{0}=PL$. Let $1\leq i,j\leq k$. We have $(KU_{0})[s_{i},C_{j}]=K_{i}U_{ij}$ and $(PL)[s_{i},C_{j}]=p_{ij}L_{j}$. So, we need to prove that $K_{i}U_{ij}=p_{ij}L_{j}$. Notice that $UA=BU$, so $\begin{pmatrix}K_{1}&\ldots&K_{k}\end{pmatrix}U_{0}=\begin{pmatrix}L_{1}&\ldots&L_{k}\end{pmatrix}$. In particular, we have that $\sum_{m=1}^{k}K_{m}U_{mj}=L_{j}$. Multiplying by $p_{ij}$ on the left and using the fact that $K_{m}$ is scalar-valued for $1\leq m\leq k$, we obtain that $p_{ij}L_{j}=\sum_{m=1}^{k}K_{m}p_{ij}U_{mj}$. Since $p_{ij}U_{mj}=\delta_{m}^{j}U_{ij}$, we reach $p_{ij}L_{j}=K_{j}U_{ij}$, as desired. This being true for all $1\leq i,j\leq k$, we have that $KU_{0}=PL$.
Now notice that $P$ and $U_{0}$ are magic unitaries, so in particular they are unitaries. So, we multiply $KU_{0}=PL$ on the left by $P^{-1}={}^{t}P$ and from the right side by $U_{0}^{-1}={}^{t}U_{0}$. We get: ${}^{t}PK={}^{t}P(KU_{0}){}^{t}U_{0}={}^{t}P(PL){}^{t}U_{0}=L{}^{t}U_{0}$. Therefore, we have ${}^{t}KP=U_{0}{}^{t}L$, as required.
This completes the proof that $V$ is a quantum isomorphism from $H^{\prime}$ to $G^{\prime}$. It remains to prove that $V$ preserves $R^{\prime}$ and $S^{\prime}$. Recall that by definition $R^{\prime}$ is the union of the blocks containing the $r_{i}$, and similarly for $S^{\prime}$. Let $x\in V(G^{\prime})$ and $a\in V(H^{\prime})$ such that $v_{ax}\neq 0$. Assume that $a\in S^{\prime}$. Then there exists $1\leq i\leq k$ such that $a$ in is the same block as $s_{i}$. We have $0\neq v_{ax}=\sum_{j=1}^{k}v_{ax}v_{s_{i}r_{j}}$, so there exists $1\leq j\leq k$ such that $v_{ax}v_{s_{i}r_{j}}\neq 0$. Since $a$ and $s_{i}$ are in a common block of $G^{\prime}$, by Lemma 3.5, we have that $x$ and $r_{j}$ are in a common block. Hence $x\in R^{\prime}$. Conversely, if $x\in R^{\prime}$, then we show in the same way that $a\in S^{\prime}$. This shows that $V$ is a quantum isomorphism of anchored graphs, as desired.
Let us now consider the case where $R$ is a block of $G$. We have that $S$ is also a block of $H$ by Lemma 4.6. Let $R^{\circ}$ (resp. $S^{\circ}$) be the vertices in $R$ (resp. in $S$) that are not cut-vertices in $G$ (resp. in $H$), so $R^{\circ}$ and $S^{\circ}$ can be empty. Using the fact that $U$ is a quantum isomorphism of anchored graphs, and Lemma 3.5, we obtain that $U$ is diagonal by blocks $U=\operatorname{Diag}(U_{1},U_{2},U_{3})$ where $U_{1}=U[S^{\circ},R^{\circ}]$, $U_{2}=U[S\setminus S^{\circ},R\setminus R^{\circ}]$, and $U_{3}=U[V(H)\setminus S,V(G)\setminus R]$. The adjacency matrix of $G$ is of the form
$A=\begin{pmatrix}A_{1}&{}^{t}K&0\\ K&A_{2}&{}^{t}L\\ 0&L&A_{3}\end{pmatrix},$
where $A_{1}$, $A_{2}$, and $A_{3}$ are respectively the adjacency matrices of $G[R^{\circ}]$, $G[R\setminus R^{\circ}]$, and $G\setminus R$. Similarly, the adjacency matrix of $H$ is of the form
$B=\begin{pmatrix}B_{1}&{}^{t}M&0\\ M&B_{2}&{}^{t}N\\ 0&N&B_{3}\end{pmatrix},$
where $B_{1}$, $B_{2}$, and $B_{3}$ are respectively the adjacency matrices of $H[S^{\circ}]$, $H[S\setminus S^{\circ}]$, and $H\setminus S$. Therefore, the adjacency matrices of $G^{\prime}$ and $H^{\prime}$ are respectively as follows:
$A^{\prime}=\operatorname{Adj}(G^{\prime})=\begin{pmatrix}0&0&0\\ 0&0&{}^{t}L\\ 0&L&A_{3}\end{pmatrix}\text{ and }B^{\prime}=\operatorname{Adj}(H^{\prime})=\begin{pmatrix}0&0&0\\ 0&0&{}^{t}N\\ 0&N&B_{3}\end{pmatrix}.$
Now, recall that $\Gamma(U)=U$. Since $UA=BU$ and $U=\operatorname{Diag}(U_{1},U_{2},U_{3})$, it is immediate to check that $UA^{\prime}=B^{\prime}U$. Let us check that $U$ preserves $R^{\prime}$ and $S^{\prime}$, that is, $u_{ax}=0$ if $x\in R^{\prime}$ and $a\notin S^{\prime}$ or $x\notin R^{\prime}$ and $a\in S^{\prime}$.
Let $a\in V(H)$ and $x\in V(G)$ such that $u_{ax}\neq 0$. First, assume that $x\in R^{\prime}$. Two cases are possible. If $x\in R$, then since $u_{ax}\neq 0$ we have by assumption that $a\in S$. Since $S\subseteq S^{\prime}$ by Lemma 4.5, we obtain that $a\in S^{\prime}$, as desired. Otherwise, $x\in R^{\prime}\setminus R$, and by Lemma 4.5 $x$ is in a common block in $G$ with a vertex $y\in R$ such that $y$ is contained in exactly one more block in $G$ other than $R$. Now, we have $0\neq u_{ax}=u_{ax}\left(\sum_{c\in V(H)}u_{cy}\right)=\sum_{c\in V(H)}u_{ax}u_{cy}.$ Hence there exists $b\in V(H)$ such that $u_{ax}u_{by}\neq 0$. Since $x$ and $y$ are in a common block in $G$, by Lemma 3.5, $a$ and $b$ are in a common block in $H$. Moreover, we have $u_{by}\neq 0$ and $y\in R$, so by assumption we obtain that $b\in S$. Since by what precedes $U$ is a quantum isomorphism from $G^{\prime}$ to $H^{\prime}$ and $y$ is not a cut vertex in $G^{\prime}$, by Lemma 3.5 $b$ is not a cut vertex in $H^{\prime}$ neither. Hence $b$ is contained in exactly one more block in $H$ than $S$. So by Lemma 4.5, we have that $a\in S^{\prime}$, as desired.
Second, applying what precedes to $U^{*}$ as a quantum isomorphism from $(H,S)$ to $(G,R)$, assuming that $u_{ax}=[U^{*}]_{xa}\neq 0$ and that $a\in S^{\prime}$, we obtain that $x\in R^{\prime}$, as desired. This shows that $U=\Gamma(U)$ is a quantum isomorphism of anchored graphs from $(G,R)$ to $(H,S)$ and concludes the proof. ∎
4.4. Operations $\Delta_{i}$
Let $(T_{1},r_{1}),\dots,(T_{k},r_{k})$ be some rooted 2-coloured trees (let us call the colours black and white). We define two operations $\Delta_{1}$ and $\Delta_{2}$, which given these rooted coloured trees, return another rooted coloured tree.
Operation $\Delta_{1}$. This operation is defined if every $r_{i}$ is white. In that case, we first consider $\bigoplus_{i=1}^{k}T_{i}$, then we add a new black vertex $r$ and join it to $r_{1},\dots,r_{k}$. The rooted 2-coloured tree $(T,r)$ obtained is $\Delta_{1}((T_{1},r_{1}),\dots,(T_{k},r_{k}))$.
Operation $\Delta_{2}$. For this operation, the color of $r_{i}$’s can be arbitrary. Among $(T_{i},r_{i})$’s, let $\mathcal{A}=\{(A_{1},a_{1}),\dots,(A_{l},a_{l})\}$ be the rooted trees whose roots are white. Also, let =$\mathcal{B}=\{(B_{1},b_{1}),\dots,(B_{m},b_{m})\}$ be the rooted trees on at least 2 vertices whose roots are black. (The rooted trees on exactly 1 vertex and whose roots are black do not play a role in the definition of $\Delta_{2}$.)
First, for each $1\leq i\leq l$, add a new new black vertex $d_{i}$ to $(A_{i},a_{i})$, join it to $a_{i}$ and set $d_{i}$ to be the new root to define a new rooted 2-coloured tree $(D_{i},d_{i})$. Then, set $T_{0}=\bigoplus_{i=1}^{l}D_{i}\oplus\bigoplus_{i=1}^{m}B_{i}$. Add a new white vertex $r$ to $T_{0}$, join it to $d_{1},\dots,d_{l},b_{1},\dots,b_{m}$ to create a 2-coloured tree. Set $r$ as the new root. We have $(T,r)=\Delta_{2}((T_{1},r_{1}),\dots,(T_{k},r_{k}))$.
Now, let us explain the connection between the operations $\Delta_{i}$ and $\Gamma$.
Lemma 4.9.
Let $(G,R)$ be a connected anchored graph. Then, $\mathcal{T}(G,R)=\Delta_{i}(\operatorname{co}(\Gamma(G,R)))$ where $\operatorname{co}(\Gamma(G,R))$ is the connected components of $\Gamma(G,R)$ (as anchored graphs) and where $i=1$ if $R$ is a cut vertex of $G$ and $i=2$ otherwise.
Proof.
The proof is immediate from the construction. ∎
5. Block structures of quantum isomorphic graphs
In this section we prove the main result of the paper. After proving some necessary lemmas, we first prove the result for anchored graphs and then deduce it for graphs.
Lemma 5.1.
Let $(T_{1},r_{1}),\dots,(T_{k},r_{k})$ and $(T^{\prime}_{1},r^{\prime}_{1}),\dots,(T^{\prime}_{k},r^{\prime}_{k})$ be rooted 2-coloured trees. Assume that for each $i\in\{1,\dots,k\}$, there exists an isomorphism $\alpha_{i}:(T_{i},r_{i})\to(T^{\prime}_{i},r^{\prime}_{i})$ (thus $\alpha_{i}$ preserves the roots and the colours).
  1. (1)
    If every $r_{i}$ is a white vertex, setting $(T,r)=\Delta_{1}((T_{1},r_{1}),\dots,(T_{k},r_{k}))$ and $(T^{\prime},r^{\prime})=\Delta_{1}((T^{\prime}_{1},r^{\prime}_{1}),\dots,(T^{\prime}_{k},r^{\prime}_{k}))$, then the extension $\alpha:(T,r)\to(T^{\prime},r^{\prime})$ of $\alpha_{i}$’s defined as below is an isomorphism of rooted 2-coloured trees:
    $\alpha(v)=\begin{cases}\alpha_{i}(v)&v\in V(T_{i})\\ r^{\prime}&v=r\end{cases}.$
  2. (2)
    Setting $(T,r)=\Delta_{2}((T_{1},r_{1}),\dots,(T_{k},r_{k}))$ and $(T^{\prime},r^{\prime})=\Delta_{2}((T^{\prime}_{1},r^{\prime}_{1}),\dots,(T^{\prime}_{k},r^{\prime}_{k}))$ and using the notations from the definition of $\Delta_{2}$, the extension $\alpha:(T,r)\to(T^{\prime},r^{\prime})$ of $\alpha_{i}$’s defined as below is an isomorphism of rooted 2-coloured trees:
    $\alpha(v)=\begin{cases}\alpha_{i}(v)&v\in V(T_{i})\\ d_{i}^{\prime}&v=d_{i}\\ r^{\prime}&v=r\end{cases}.$
Proof.
(1) is immediate from the definition of $\Delta_{1}$.
For (2), we need to verify that for $i\in\{1,\dots,k\}$, if $d_{i}x\in E(T)$, then $d^{\prime}_{i}\alpha(x)\in E(T^{\prime})$ and the rest is immediate from the definition of $\Delta_{2}$. So, let $x\in V(T)$ such that $d_{i}x\in E(T)$. Therefore, using the notation in the definition of $\Delta_{2}$, either $x=r$ or $x=a_{i}$ for some $i\in\{1,\dots,l\}$. In the former case, $\alpha(x)=r^{\prime}$ and thus by construction of $T^{\prime}$, we have $d^{\prime}_{i}\alpha(x)\in E(T^{\prime})$. In the latter case, notice that $\alpha(a_{i})=\alpha_{i}(a_{i})=a^{\prime}_{i}$ (because $\alpha_{i}$ preserves the roots) and therefore, $d_{i}^{\prime}\alpha(a_{i})\in E(T^{\prime})$. This concludes the proof. ∎
Theorem 5.2.
Let $(G,R)$ and $(H,S)$ be two quantum isomorphic connected anchored graphs. Then:
  1. (1)
    there exists $\alpha:\mathcal{T}(G,R)\rightarrow\mathcal{T}(H,S)$ such that
    • $\alpha$ is an isomorphism between the block trees of $G$ and $H$ preserving the roots,
    • for every block $b$ of $G$, we have $b\simeq_{q}\alpha(b)$.
  2. (2)
    there exists $\beta:\mathcal{B}(G,R)\rightarrow\mathcal{B}(H,S)$ such that
    • $\beta$ is an isomorphism between the block graphs of $G$ and $H$ preserving the anchors,
    • for every block $b$ of $G$, we have $b\simeq_{q}\beta(b)$.
Proof.
Let us start by proving Item (1) by induction on the number of blocks of $G$.
Set $(G^{\prime},R^{\prime})=\Gamma(G,R)$ and $(H^{\prime},S^{\prime})=\Gamma(H,S)$. By Lemma 4.8, $V=\Gamma(U)$ is a quantum isomorphism from $(G^{\prime},R^{\prime})$ to $(H^{\prime},S^{\prime})$. Remember that these two anchored graphs are not connected. Using the same notation as in the definition of $\Gamma$, we have $(G^{\prime},R^{\prime})=\bigoplus_{i=1}^{k}(G_{i},R_{i})$. Similarly, $(H^{\prime},S^{\prime})=\bigoplus_{i=1}^{k^{\prime}}(H_{i},S_{i})$. By Lemma 4.2, up to reordering, we have that $k=k^{\prime}$ and that $(G_{i},R_{i})$ is quantum isomorphic to $(H_{i},S_{i})$. By induction hypothesis, for every $i\in\{1,\dots,k\}$ there exists an isomorphism of rooted trees $\alpha_{i}:\mathcal{T}(G_{i},R_{i})\to\mathcal{T}(H_{i},S_{i})$ such that for every block $b$ of $G_{i}$, we have $b\simeq_{q}\alpha_{i}(b)$.
Recall that one can see $\mathcal{T}(G_{i},R_{i})$ and $\mathcal{T}(H_{i},S_{i})$ as 2-coloured graphs by assigning the colour black to every vertex that is a cut vertex of the original graph and the colour white to every other vertex. Remember that if the tree has only one vertex, this vertex is coloured black in our setting. Notice that by Lemma 4.3 and the fact that $\alpha_{i}$ preserves the roots, we have that the roots $R_{i}$ and $S_{i}$ have the same color. Moreover, since the color of a vertex is uniquely determined by its distance from the root, we deduce that $\alpha_{i}$ preserves the colours. Thus each $\alpha_{i}$ is an isomorphism between rooted 2-coloured trees.
So, by Lemma 5.1, we can find an isomorphism $\alpha^{(j)}:\Delta_{j}(\mathcal{T}(G_{1},R_{1}),\dots,\mathcal{T}(G_{k},R_{k}))\to\Delta_{j}(\mathcal{T}(H_{1},S_{1}),\dots,\mathcal{T}(H_{k},S_{k}))$ for $j=1$ when $\Delta_{1}$ is defined and for $j=2$, in any case.
Now, if $R=\{r\}$ where $r$ is a cut vertex of $G$, then by Lemma 3.5 $S$ is also a cut vertex if $G$. So by Lemma 4.9, $\mathcal{T}(G,R)=\Delta_{1}(\mathcal{T}(G_{1},R_{1}),\dots,\mathcal{T}(G_{k},R_{k}))$ and $\mathcal{T}(H,S)=\Delta_{1}(\mathcal{T}(H_{1},S_{1}),\dots,\mathcal{T}(H_{k},S_{k}))$. Thus $\alpha=\alpha^{(1)}$ is the desired isomorphism. Moreover, for every block $b$ of $G$, we have that $b$ is a white vertex in $\mathcal{T}(G,R)$, thus is particular $b\neq R$ (as vertices of $\mathcal{T}(G,R)$). Therefore, $\alpha(b)=\alpha_{i}(b)$ (for some $i$) which, by induction hypothesis, is quantum isomorphic to $b$. Thus, $\alpha(b)\simeq_{q}b$.
If $R$ is not a cut vertex of $G$, then $S$ is not a cut vertex of $H$. Moreover, by Lemma 4.9, $\mathcal{T}(G,R)=\Delta_{2}(\mathcal{T}(G_{1},R_{1}),\dots,\mathcal{T}(G_{k},R_{k}))$ and $\mathcal{T}(H,S)=\Delta_{2}(\mathcal{T}(H_{1},S_{1}),\dots,\mathcal{T}(H_{k},S_{k}))$. In this case, $\alpha=\alpha^{(2)}$ is the desired isomorphism. Moreover, if $b$ is a block of $G$, then either $b=R$, in which case $\alpha(b)=S$ and by Lemma 4.6, $b\simeq_{q}\alpha(b)$. Or $b$ is a block of $G_{i}$ for some $i$. Thus by induction hypothesis $\alpha(b)=\alpha_{i}(b)\simeq_{q}b$.
To prove (2), we define $\beta:\mathcal{B}(G,R)\rightarrow\mathcal{B}(H,S)$ using $\alpha$ from Item (1). For $b\in V(\mathcal{B}(G,R))$, we define $\beta(b)=\alpha(b)$: this makes sense since $b$ is a block of $G$, hence a (white) vertex of $\mathcal{T}(G,R)$.
Let us first verify that $\beta$ is a graph isomorphism from $\mathcal{B}(G)$ to $\mathcal{B}(H)$. Since $\alpha$ induces a bijection between the blocks of $G$ and the blocks of $H$, we already know $\beta$ is a bijection. Let us check it is a morphism first. Let $bb^{\prime}\in E(\mathcal{B}(G))$. This means that there exists a cut vertex $c$ of $G$ such that $c$ is in the intersection of $b$ and $b^{\prime}$. Since $c\in b$ and $c\in b^{\prime}$, we have that $cb\in E(\mathcal{T}(G))$ and $cb^{\prime}\in E(\mathcal{T}(G))$. Therefore, $\alpha(c)\alpha(b),\alpha(c)\alpha(b^{\prime})\in E(\mathcal{T}(H))$. This implies that $\alpha(c)\in\alpha(b)$ and $\alpha(c)\in\alpha(b^{\prime})$. Moreover, we know that $\alpha$ preserves colours. Thus, $\alpha(c)$ is a cut vertex of $H$ in the intersection of the blocks $\alpha(b)$ and $\alpha(b^{\prime})$ of $H$. Hence $\beta(b)\beta(b^{\prime})=\alpha(b)\alpha(b^{\prime})\in E(\mathcal{B}(H))$, so $\beta$ is a graph morphism.
To prove that $\beta$ is an isomorphism, it remains to check the preimage of an edge is an edge. For this, assume that $\beta(b)\beta(b^{\prime})\in E(\mathcal{B}(H))$ for some $b$, $b^{\prime}\in V(\mathcal{B}(G))$. In particular, there exists a cut vertex $x\in V(H)$ such that $x\in\beta(b)\cap\beta(b^{\prime})$. Since $\alpha$ is surjective, there is a cut vertex $c\in V(G)$ such that $x=\alpha(c)$. Moreover, $\alpha$ is a graph isomorphism, so $cb\in E(\mathcal{T}(G))$ and $cb^{\prime}\in E(\mathcal{T}(G))$. Hence $c\in b\cap b^{\prime}$ and $bb^{\prime}\in E(\mathcal{B}(G))$, as desired. This shows that $\beta$ is a graph isomorphism.
Now, let us show that $\beta$ preserves the anchors. Let $\mathcal{B}(G,R)=(\mathcal{B}(G),\tilde{R})$ and $\mathcal{B}(H,S)=(\mathcal{B}(H),\tilde{S})$, we want to prove that $\beta(\tilde{R})=\tilde{S}$. If $R$ is a cut vertex of $G$, then $S$ is a cut vertex of $H$. Moreover, in this case, $\tilde{R}=\nu_{G}(R)=N_{\mathcal{T}(G)}(R)$ and $\tilde{S}=\nu_{H}(S)=N_{\mathcal{T}}(H)(S)$. Since $\alpha$ is an isomorphism sending $R$ to $S$, it also sends the neighbourhood of $R$ to the neighbourhood of $S$. Thus
$\beta(\tilde{R})=\alpha(N_{\mathcal{T}(G)}(R))=N_{\mathcal{T}(H)}(\alpha(R))=N_{\mathcal{T}(H)}(S)=\tilde{S},$
as desired.
If $R$ is a block of $G$, then $S$ is a block of $H$ by Lemma 4.6, and $\tilde{R}=\rho_{T(G)}(R)$ and $\tilde{S}=\rho_{T(H)}(S)$. We consider two cases. First, if $R$ is an internal vertex in $\mathcal{B}(G)$, then since $\beta$ is an isomorphism, $S$ is also an internal vertex of $\mathcal{B}(H)$. Moreover, in this case, $\tilde{R}=\rho_{\mathcal{T}(G)}(R)=N_{\mathcal{B}(G)}[R]$ and $\tilde{S}=\rho_{\mathcal{T}(H)}(S)=N_{\mathcal{B}(H)}[S]$. Since $\beta$ is an isomorphism, it sends the closed neighbourhood of a vertex $x$ to the closed neighbourhood of $\beta(x)$. Thus,
$\beta(\tilde{R})=\alpha(N_{\mathcal{B}(G)}[R])=N_{\mathcal{B}(H)}[\alpha(R)]=N_{\mathcal{B}(H)}[S]=\tilde{S},$
as desired.
Second, if $R$ is a cut vertex in $\mathcal{B}(G)$, then $S$ is a cut vertex in $\mathcal{B}(H)$. So,
$\beta(\tilde{R})=\beta(\rho_{\mathcal{T}(G)}(R))=\beta(\{R\})=\{\beta(R)\}=\{\alpha(R)\}=\{S\}=\rho_{\mathcal{T}(H)}(S)=\tilde{S},$
as desired.
Finally, for $b$ a block of $G$, we know that $b\simeq_{q}\alpha(b)=\beta(b)$. This completes the proof of the theorem. ∎
Now we can prove the main theorem of this section.
Theorem 5.3.
If $G$ and $H$ are two quantum isomorphic graphs, then
  1. (1)
    there exists $\alpha:\mathcal{T}(G)\rightarrow\mathcal{T}(H)$ such that
    • $\alpha$ is an isomorphism between the block trees of $G$ and $H$,
    • for every block $b$ of $G$, we have $b\simeq_{q}\alpha(b)$.
  2. (2)
    there exists $\beta:\mathcal{B}(G)\rightarrow\mathcal{B}(H)$ such that
    • $\beta$ is an isomorphism between the block graphs of $G$ and $H$,
    • for every block $b$ of $G$, we have $b\simeq_{q}\beta(b)$.
Proof.
By Lemma 3.4, it is enough to prove the result for connected graphs. Hence we assume both $G$ and $H$ to be connected. Let $R=\rho_{G}(Z(G))$ and $S=\rho_{H}(Z(H))$ and consider the two anchored graphs $(G,R)$ and $(H,S)$. By Lemma 3.7, these two anchored graphs are quantum isomorphic. Moreover, notice that the block graphs and the block tree of $G$ and $H$ are the same as the underlying block graphs and block trees of $(G,R)$ and $(H,S)$. Therefore, we conclude by Theorem 5.2. ∎
Corollary 5.4.
Let $(G,H)$ be a minimal pair of quantum isomorphic graphs which are not isomorphic. Then $G$ and $H$ are 2-connected.
Finally, we obtain that the number of blocks containing a vertex is preserved under quantum isomorphism.
Theorem 5.5.
Let $U$ be a quantum isomorphism from $G$ to $H$. Let $x\in V(G)$ and $a\in V(H)$. If $u_{ax}\neq 0$, then the number of blocks of $G$ containing $x$ is equal to the number of blocks of $H$ containing $a$.
Proof.
Let us denote the number of blocks containing a vertex $v$ in a graph $F$ by $D_{F}(v)$.
If $x$ is an internal vertex of $G$, then $a$ is also an internal vertex of $H$ by Lemma 3.5, and the result is immediate since $D_{G}(x)=1=D_{H}(a)$. So, for the rest of the proof, we assume that $x$ is a cut vertex (thus so is $a$).
We first prove the following statement for anchored graphs: let $U$ be a quantum isomorphism from an anchored graph $(G,R)$ to an anchored graph $(H,S)$ and let $x\in G$ and $a\in H$ be cut vertices. If $u_{ax}\neq 0$, then $D_{G}(x)=D_{H}(a)$.
We prove the statement by induction on the number of blocks of $G$ (which, by Theorem 5.3, is equal to the number of blocks of $H$).
First, assume that $R=\{r\}$ where $r$ is a cut vertex of $G$. So by Lemma 3.5 $S=\{s\}$ with $s$ a cut vertex of $H$, and $U=\operatorname{Diag}[U_{0},1]$. Thus either $(x,a)=(r,s)$ or we have $x\neq r$ and $a\neq s$. Since $G\setminus R$ is quantum isomorphic to $H\setminus S$ through $U_{0}$, the number of their connected components is equal by Lemma 3.4. But the number of connected components of $G\setminus R$ (resp. $H\setminus S$) is exactly $D_{G}(r)$ (resp. $D_{H}(s)$). Thus $D_{G}(r)=D_{H}(s)$. So, we only need to consider the case where $x\neq r$, $a\neq s$, and $u_{ax}\neq 0$. Notice that in this case, $x$ and $a$ are cut vertices in $(G^{\prime},R^{\prime})=\Gamma(G,R)$ and $(H^{\prime},S^{\prime})=\Gamma(H,S)$ respectively. Moreover, setting $V=\Gamma(U)=\operatorname{Diag}[U_{0},P(U_{0})]$, we have by Lemma 4.8 that $V$ is a quantum isomorphism from $\Gamma(G,R)$ to $\Gamma(H,S)$. Since $x\neq r$ and $a\neq s$, we have $x\in V(G^{\prime})$ and $a\in V(H^{\prime})$ and $v_{ax}=u_{ax}\neq 0$. Moreover, by Lemma 4.2, the anchored connected components of $(G,R)$ and $(H,S)$ are two-by-two quantum isomorphic, and they have strictly less blocks than $G$. So we can apply the induction hypothesis to each pair of quantum isomorphic anchored connected components. Using the fact that in the current case the number of blocks containing $x$ and containing $a$ remains the same under the operation $\Gamma$, we have that
$D_{G}(x)=D_{\Gamma(G,R)}(x)=D_{\Gamma(H,S)}(a)=D_{H}(a),$
as desired.
Second, assume that $R$ is not a cut vertex of $G$. Therefore, $S$ is not a cut vertex of $H$ neither. Let $(G^{\prime},R^{\prime})=\Gamma(G,R)$ and $(H^{\prime},S^{\prime})=\Gamma(H,S)$. Notice that $U=\Gamma(U)$ is also a quantum isomorphism from $(G^{\prime},R^{\prime})$ to $(H^{\prime},S^{\prime})$ by Lemma 4.8. So, if $x$ is no more a cut vertex in $G^{\prime}$, then $a$ is no more a cut vertex in $H^{\prime}$ since $u_{ax}\neq 0$ and by Lemma 3.5. But in that case, $x$ (resp. $a$) must have been in exactly one block other than $R$ (resp. $S$). Therefore, $D_{G}(x)=2=D_{H}(a)$.
On the other hand, if $x$ remains a cut vertex in $G^{\prime}$, then $a$ also remains a cut vertex in $H^{\prime}$. Now, two cases are possible: either $x\in R$ (and thus $a\in S$) or $x\notin R$ (thus $a\notin S$). In the latter case, by induction hypothesis, we have that
$D_{G}(x)=D_{\Gamma(G,R)}(x)=D_{\Gamma(H,S)}(a)=D_{H}(a).$
In the former case, again by induction hypothesis, we have:
$D_{G}(x)=D_{\Gamma(G,R)}(x)+1=D_{\Gamma(H,S)}(a)+1=D_{H}(a).$
This completes the proof for anchored graphs.
Now, let us prove the statement of the theorem for graphs. Set $R=\rho_{G}(Z(G))$ and $S=\rho_{H}(Z(H))$. By Lemma 3.7, these two anchored graphs are quantum isomorphic. So, the result follows from what precedes. ∎
6. Open problems
Let us now discuss a few open problems that arise naturally from the results of last section.
We say that a connected graph $G$ is $k$-connected if it has at least $k$ vertices and for every $S\subseteq V(G)$ of size at most $k-1$, the graph $G\setminus S$ is connected. The connectivity of a connected graph $G$ is the maximum $k\in\mathbb{N}$ such that $G$ is $k$-connected.
Question 6.1.
Does there exist a pair of quantum isomorphic graphs $G$ and $H$ whose connectivities are not the same?
Two graphs are cospectral if the spectra of their adjacency matrices are the same. It is well-known that quantum isomorphic graphs are cospectral. Indeed, since a magic unitary is a unitary, the adjacency matrices of the graphs are then conjugate by a unitary in a $C^{*}$-algebra, and since the spectrum of an element in a $C^{*}$-algebra does not change when going to a bigger $C^{*}$-algebra, we obtain that the adjacency matrices have the same spectrum.
If the answer to 6.1 is positive, then it can be seen as a generalisation of the following theorem.
Theorem 6.2.
There exist cospectral graphs with different connectivities.
For examples of such pairs, see [Hae19, Theorem 2.1] where for every $k\geq 2$ it is shown that there exist $k$-regular cospectral graphs such that the connectivity of one is $2k$ and the connectivity of the other is $k+1$. However, we remark that cospectrality preserves being connected.
Moreover, if the answer to 6.1 is positive, one can further ask the following question.
Question 6.3.
What is the maximum $k$ such that $k$-connectivity is preserved under quantum isomorphism? That is, what is the maximum $k$ such that for every $k$-connected graph $G$ we have that every graph quantum isomorphic to $G$ is $k$-connected?
By Theorem 5.3, we have that $k\geq 2$ and if the answer to 6.1 is positive, then $k<+\infty$.
Let $(G,H)$ be a pair of quantum isomorphic graphs that are not isomorphic. We say that it is an irreducible such pair if for every induced subgraph $G^{\prime}$ of $G$ and every induced subgraph $H^{\prime}$ of $H$, if $G^{\prime}$ and $H^{\prime}$ are quantum isomorphic, then they are isomorphic.
As described in Corollary 5.4, if $(G,H)$ is an irreducible such pair, then both graphs should be 2-connected. The following question asks whether one can improve this bound.
Question 6.4.
Does there exist an irreducible pair $(G,H)$ of quantum isomorphic graphs that are not quantum isomorphic and such that $G$ is not 3-connected?
Let us conclude by mentioning the following question.
Question 6.5.
Are there infinitely many irreducible pairs of quantum isomorphic graphs that are not isomorphic?